# The Belle II Collaboration

### v1

@@ III Event reconstruction and selection
The inclusive $B\rightarrow X_{u}\ell\nu$ analysis relies on the reconstruction of the $B_{\mathrm{tag}}$ meson in hadronic decay channels. The rest frame of the signal $B$ meson is determined from momentum conservation in $\Upsilon(4S)\to B\bar{B}$ decays. The lepton in the semileptonic decay is reconstructed as a single track. All detected tracks and clusters not associated with the $B_{\mathrm{tag}}$ or the lepton are assumed to come from the $X$ system and are grouped into an object called the rest-of-event (ROE). The hadronic system four-momentum $p_{X}$, from which the hadronic mass $M_{X}$ is derived, is computed from the four-momenta of all tracks and clusters in the ROE. All ROE clusters are assumed to be produced by photons and, since events with kaons and leptons in the ROE are rejected at a later stage, all ROE tracks are assigned a pion mass hypothesis. Particles originating from beam background can pollute the ROE and they are therefore suppressed by the selections given in Section III.2.2. The neutrino cannot be detected and its momentum is therefore identified with the missing momentum of the event. Using the four momenta of the initial $e^{+}e^{-}$ state, the $B_{\mathrm{tag}}$, the lepton and $X$, the neutrino four-momentum can be inferred from the missing four-momentum as:
@@
In the following, the requirements applied on $B_{\mathrm{tag}}$ candidates, ROE particles and signal lepton candidates are presented. The requirements described in Sections III.1 and III.2 are also referred to as preselections.
@@ III.2.2 Rest-Of-Event
ROE selections are optimized to select charged particles originating near the interaction point and suppress particles that hit the edges of the detector where the response is poorly modelled. Selections applied on neutral clusters are optimized to suppress beam background and misidentified photons. A fit is performed using all tracks found in the ROE to match them to a single decay vertex. Tracks in the ROE are required to pass similar $dr$, dzdz and polar angle selections to signal lepton tracks. Duplicated tracks due to particles with low transverse momentum are suppressed by a BDT. The classifier uses the track helix parameters, the product of the track charges, the angle between the tracks and the track transverse and longitudinal momenta.
@@ III.2.3 Charged pions
The multiplicity of charged pions in an event is used to derive an uncertainty on the $X_{u}$ hadronization modeling as described in Section V.5.13. These pions are identified using a likelihood ratio score calculated from information collected by all subdetectors except the PXD and are required to have a momentum and polar angle within the phase-space region considered in hadron identification calibration studies [36]. The likelihood requirement selects about 47% of the charged pion candidates and about 2% of pion candidates are misidentified particles.
@@ III.3 Continuum
In $e^{+}e^{-}\to q\bar{q}$ events, light mesons are boosted and therefore result in two back-to-back jet-like structures, whereas $B$ mesons are produced nearly at rest in the center-of-mass frame resulting in isotropically distributed decay products. Variables such as the two $B$ meson thrust angles and magnitudes, the modified Fox-Wolfram moments [37, 38] and the CLEO cones [39] can be used to distinguish continuum events from $B$ meson decays (30 variables are considered). In order to suppress the continuum background, a multi layer perceptron (MLP) is trained. The MLP is configured with two hidden layers with 256 and 128 neurons, respectively. The network uses the Swish activation function [40] and a binary cross-entropy loss function. The AdamW optimizer [41] is used to train the network, with a learning rate of $9.3\times 10^{-3}$. The classifier is trained using a total of 14 variables: the charged track multiplicity, and a subset of 13 variables among the 30 describing the event shape as mentioned above. [^3] The other 17 variables are dropped because they display poor agreement between experimental and simulated data in the region used to train the classifier. All chosen variables exhibit good agreement between simulated and observed data. A subset consisting of 5% of simulated $B\rightarrow X_{u}\ell\nu$ events and 20% of simulated continuum events is used for training. The sample of continuum events is corrected following the procedure described in Section IV.1. The neural network is optimized using the optuna package [42]. The optimal selection on the classifier output score is chosen according to the maximal significance defined as $S/\sqrt{S+B}$, with $S$ the number of $B\rightarrow X_{u}\ell\nu$ events and $B$ the number of continuum events that pass the selection. The yields are counted in the region where $E_{\ell}^{B}>1$ GeV (see Section III.6 for a discussion on kinematic selections). The continuum suppression selection rejects 95% of continuum events while retaining 68% of $B\rightarrow X_{u}\ell\nu$ events.
@@ III.5 $B\rightarrow X_{c}\ell\nu$ suppression
Since the average particle multiplicity of $B\rightarrow X_{c}\ell\nu$ events is expected to be larger than for $B\rightarrow X_{u}\ell\nu$ events, the event reconstruction quality is expected to be poorer because of missed particles or because of the limited detector resolution. The first three variables therefore help to reject poorly reconstructed $B\rightarrow X_{c}\ell\nu$ events which tend to populate the high $M_{\rm{miss}}^{2}$ tail, have a low $p_{\rm vtx}^{\rm ROE}$ and a $Q_{\rm{tot}}$ different than zero. The $\pi_{s}$ variables allow the identification of events with a $D^{*}$ decay. The distributions of the nine input features are shown in Appendix A. The optimal selection on the classifier output score is chosen as the one which minimizes the $\Delta\mathcal{B}(B\rightarrow X_{u}\ell\nu)$ total uncertainty as obtained from the nominal fit described in Section V. The output score is shown in Figure 2 and the optimal selection value for the score is found to be 0.87. The $B\rightarrow X_{c}\ell\nu$ suppression requirement rejects 98% of $B\rightarrow X_{c}\ell\nu$ events while retaining 25% of $B\rightarrow X_{u}\ell\nu$ events.
Figure 2: $B\rightarrow X_{c}\ell\nu$ suppression classifier output score. Simulated templates are shown as stacked histograms and experimental data are shown as black points. The signal component is shown as a purple line with its yield multiplied by 10 to enhance visibility. The bottom panel shows the ratio between data and simulation yields in each bin. The purple arrows indicate that the point is outside of the plotting range and the purple crosses indicate the absence of simulated data in a particular bin. The step in the distribution in the range 0.700.750.70-0.75 is caused by the large number of events with a $p_{\rm vtx}^{\rm ROE}$ close to zero (see Figure 10 in Appendix A).
@@ IV.1 Continuum modeling corrections
The modeling of continuum events is known to be poorly understood and is therefore corrected using a data-driven approach. First, the total number of expected events is corrected by comparing the number of events in 40% of the off-resonance experimental and simulated data samples. The remaining 60% are used to test and validate the BDT described below. The correction factor is extracted independently for events in which a charged or neutral $B$ meson candidate has been reconstructed and is found to be $0.87\pm 0.02$ and $0.94\pm 0.03$, respectively, where the uncertainties are statistical (uncertainties on these factors are derived as described in Section V.5.9).
In a second step, a BDT with a maximum tree depth of five is trained to distinguish between experimental and simulated off-resonance data using 30 variables that describe the event shape [43] (see Section III.3). In addition, the missing mass squared and the scalar sum of the transverse momenta of the signal and tag-side $B$ final-state particles are included in the BDT input features. The BDT is trained on the off-resonance sample from which the normalization correction factors are extracted. The remainder of the sample is split into a validation sample (10% of the full sample) and a test sample (50%). The classifier output score $\mathcal{S}_{\rm cont}$ is well calibrated and takes values between 0 and 1. The ratio
@@
is an approximation to the ratio of experimental to simulated data likelihoods: it is applied as a weight to continuum events in the on-resonance simulated data sample. The continuum correction includes both the overall normalization factors and the BDT-extracted shape factors. The modeling correction is applied on the continuum sample after applying the preselections and before training the continuum suppression classifier. Its impact is illustrated in Figure 3 for the most discriminating variable in the BDT training.
Figure 3: Distribution of $C_{1}$ (first CLEO cone [39]) in an off-resonance sample before (left) and after (right) the continuum modeling correction is applied. The modeling uncertainties, described in Section V.5.9, are included in the error band after applying the continuum calibration. The bottom panel shows the ratio between data and simulation yields in each bin. The purple arrows indicate that the point is outside of the plotting range and the purple crosses indicate the absence of simulated data in a particular bin.
@@ IV.2 $B\rightarrow X_{c}\ell\nu$ modeling corrections
The agreement between data and simulation is evaluated in control regions. Three control regions are defined for each kinematic selection defined in Table 2 by inverting the kaon veto and/or requiring a low $B\rightarrow X_{c}\ell\nu$ suppression classifier score (see Figure 2):
@@
A comparison of the $B\rightarrow X_{c}\ell\nu$ composition in the three control regions and the signal region is shown in Appendix B. In addition, two validation regions, VRK,mid and VR0,mid, with classifier scores between 0.60 and 0.87 are defined. In the former region, a kaon is required whereas in the latter, kaons are vetoed. The subdivided selection plane is illustrated in Figure 4. In Figure 5, clear disagreements between data and simulation distributions of $E_{\ell}^{B}$, $M_{X}$ and $q^{2}$ can be seen.
@@
Background originating from continuum as well as fake and secondary leptons is known to be poorly modelled in simulation. However, continuum events are corrected following the procedure described in Section IV.1 and these background sources represent in total less than 10% of all events in the control regions. The mismodeling is therefore expected to stem from $B\rightarrow X_{c}\ell\nu$ events. There are several known shortcomings in the simulation of $B\rightarrow X_{c}\ell\nu$ decays. The $B\to D^{(*)}\eta\ell\nu$ modes have never been observed and are used as a hypothesis for the decays that fill the $B\rightarrow X_{c}\ell\nu$ gap. Therefore, a relative uncertainty of 100% is assigned to their assumed branching fractions. This uncertainty doesn’t cover the full difference between data and simulation distributions shown in Figure 5. Furthermore, most $D$ meson decays with three or more particles are not simulated according to measured decay distributions but use simple phase-space models, which can bias the kinematics of charmed $B$ decays. Moreover, it was observed in Belle II studies [44] that measurements can be affected by the poor modeling of charm decays involving $K^{0}_{L}$ mesons. The impact of the modeling of $B\rightarrow X_{c}\ell\nu$ events with a $D$ meson decaying to a $K^{0}_{L}$ is studied in detail by scaling their contribution by arbitrary factors. Based on studies performed in the measurement of the $B^{+}\to K^{+}\nu\bar{\nu}$ branching fraction [44], this component is increased by 30% which only marginally improves the agreement between data and simulation. This indicates that it is unlikely that the mismodeling of $D\to K^{0}_{L}X$ decays can explain the observed mismodeling. As no single source of mismodeling could be identified, the $B\rightarrow X_{c}\ell\nu$ component is corrected as a whole using the control regions defined above. The normalization and shape from simulation are corrected separately.
Figure 5: $E_{\ell}^{B}$, $M_{X}$ and $q^{2}$ distributions in data and simulation in the CR0,low region defined in the text. Only the $E_{\ell}^{B}>1$ GeV kinematic selection is applied. All systematic uncertainties as described in Section V.5 are included. The bottom panel shows the ratio between data and simulation yields in each bin. The purple arrows indicate that the point is outside of the plotting range.
@@ IV.3 $B\rightarrow X_{c}\ell\nu$ normalization correction
The $B\rightarrow X_{c}\ell\nu$ normalization correction factors $r_{0,\mathrm{low}}$, $r_{K,\mathrm{low}}$ and $r_{K,\mathrm{high}}$ are calculated by subtracting from data and simulation all $B\rightarrow X_{u}\ell\nu$, continuum, as well as fake and secondary lepton simulated events. These components represent between 2% and 10% of all events in CR0,low and CRK,low and between 8% and 25% of events in CRK,high, depending on kinematic selections. Since this computation relies on various sources of background and signal being well modelled, an uncertainty is derived on the three ratios by varying the total yields of $B\rightarrow X_{u}\ell\nu$, continuum and fake and secondary lepton events by 50% and computing the normalization correction factors for each variation. The calculation defined in Equation 10 is tested by replacing CRK,high and the signal region by VRK,mid and VR0,mid  respectively. Although the signal purity is larger in VR0,mid (about 10–20% depending on kinematic selections) than in all other control regions, it remains lower than in the signal region (30–70%). The $B\rightarrow X_{c}\ell\nu$ normalization correction obtained for VR0,mid via the control region method (1.07, 1.15, 1.21, from the broadest set of kinematic selections to the tightest) and the factor obtained by directly taking the $B\rightarrow X_{c}\ell\nu$ data over simulated data yield ratio in that region (1.03, 1.16, 1.22) are compared. The relative difference between these two factors is considered as an additional uncertainty on $r_{\rm SR}$. In the signal extraction, two nuisance parameters are included to float the normalization of the $B\rightarrow X_{c}\ell\nu$ component in CR0,low and in the signal region separately. The normalization parameters are constrained by Gaussian functions with widths equal to the uncertainty on the normalization correction of each region. The values of $r_{0,\mathrm{low}}$ and $r_{\rm SR}$ in the three phase-space regions considered in the signal extraction are given in Table 3.
@@
The $B\rightarrow X_{c}\ell\nu$ shape correction is implemented directly in the signal extraction procedure by simultaneously fitting the signal region and CR0,low as described in Section V.2.

### v2

@@ III Event reconstruction and selection
The inclusive $B\rightarrow X_{u}\ell\nu$ analysis relies on the reconstruction of the $B_{\mathrm{tag}}$ meson in hadronic decay channels. The rest frame of the signal $B$ meson is determined from momentum conservation in $\Upsilon(4S)\to B\bar{B}$ decays. The lepton in the semileptonic decay is reconstructed as a single track. All detected tracks and clusters not associated with the $B_{\mathrm{tag}}$ or the lepton are assumed to come from the $X$ system and are grouped into an object called the rest-of-event (ROE). The hadronic system four-momentum $p_{X}$, from which the hadronic mass $M_{X}$ is derived, is computed from the four-momenta of all tracks and clusters in the ROE. All ROE clusters are assumed to be produced by photons and, since events with kaons and leptons in the ROE are rejected at a later stage, all ROE tracks are assigned a pion mass hypothesis. Particles originating from beam background can pollute the ROE and they are therefore suppressed by the selections given in Sec. III.2.2. The neutrino cannot be detected and its momentum is therefore identified with the missing momentum of the event. Using the four momenta of the initial $e^{+}e^{-}$ state, the $B_{\mathrm{tag}}$, the lepton and $X$, the neutrino four-momentum can be inferred from the missing four-momentum as:
@@
In the following, the requirements applied on $B_{\mathrm{tag}}$ candidates, ROE particles and signal lepton candidates are presented. The requirements described in Secs III.1 and III.2 are also referred to as preselections.
@@ III.2.2 Rest-Of-Event
ROE selections are optimized to select charged particles originating near the interaction point and suppress particles that hit the edges of the detector where the response is poorly modeled. Selections applied on neutral clusters are optimized to suppress beam background and misidentified photons. A fit is performed using all tracks found in the ROE to match them to a single decay vertex. Tracks in the ROE are required to pass similar $dr$, dzdz and polar angle selections to signal lepton tracks. Duplicated tracks due to particles with low transverse momentum are suppressed by a BDT. The classifier uses the track helix parameters, the product of the track charges, the angle between the tracks and the track transverse and longitudinal momenta.
@@ III.2.3 Charged pions
The multiplicity of charged pions in an event is used to derive an uncertainty on the $X_{u}$ hadronization modeling as described in Sec. V.5.13. These pions are identified using a likelihood ratio score calculated from information collected by all subdetectors except the PXD and are required to have a momentum and polar angle within the phase-space region considered in hadron identification calibration studies [3]. The likelihood requirement selects about 47% of the charged pion candidates and about 2% of pion candidates are misidentified particles.
@@ III.3 Continuum
In $e^{+}e^{-}\to q\bar{q}$ events, light mesons are boosted and therefore result in two back-to-back jetlike structures, whereas $B$ mesons are produced nearly at rest in the center-of-mass frame resulting in isotropically distributed decay products. Variables such as the two $B$ meson thrust angles and magnitudes, the modified Fox-Wolfram moments [34, 44] and the CLEO cones [11] can be used to distinguish continuum events from $B$ meson decays (30 variables are considered). In order to suppress the continuum background, a multi layer perceptron (MLP) is trained. The MLP is configured with two hidden layers with 256 and 128 neurons, respectively. The network uses the Swish activation function [55] and a binary cross-entropy loss function. The AdamW optimizer [47] is used to train the network, with a learning rate of $9.3\times 10^{-3}$. The classifier is trained using a total of 14 variables: the charged track multiplicity, and a subset of 13 variables among the 30 describing the event shape as mentioned above. [^3] The other 17 variables are dropped because they display poor agreement between experimental and simulated data in the region used to train the classifier. All chosen variables exhibit good agreement between simulated and observed data. A subset consisting of 5% of simulated $B\rightarrow X_{u}\ell\nu$ events and 20% of simulated continuum events is used for training. The sample of continuum events is corrected following the procedure described in Sec. IV.1. The neural network is optimized using the OPTUNA package [7]. The optimal selection on the classifier output score is chosen according to the maximal significance defined as $S/\sqrt{S+B}$, with $S$ the number of $B\rightarrow X_{u}\ell\nu$ events and $B$ the number of continuum events that pass the selection. The yields are counted in the region where $E_{\ell}^{B}>1$ GeV (see Sec. III.6 for a discussion on kinematic selections). The continuum suppression selection rejects 95% of continuum events while retaining 68% of $B\rightarrow X_{u}\ell\nu$ events.
@@ III.5 $B\rightarrow X_{c}\ell\nu$ suppression
Since the average particle multiplicity of $B\rightarrow X_{c}\ell\nu$ events is expected to be larger than for $B\rightarrow X_{u}\ell\nu$ events, the event reconstruction quality is expected to be poorer because of missed particles or because of the limited detector resolution. The first three variables therefore help to reject poorly reconstructed $B\rightarrow X_{c}\ell\nu$ events which tend to populate the high $M_{\rm{miss}}^{2}$ tail, have a low $p_{\rm vtx}^{\rm ROE}$ and a $Q_{\rm{tot}}$ different than zero. The $\pi_{s}$ variables allow the identification of events with a $D^{*}$ decay. The distributions of the nine input features are shown in Appendix A. The optimal selection on the classifier output score is chosen as the one which minimizes the $\Delta\mathcal{B}(B\rightarrow X_{u}\ell\nu)$ total uncertainty as obtained from the nominal fit described in Sec. V. The output score is shown in Fig. 2 and the optimal selection value for the score is found to be 0.87. The $B\rightarrow X_{c}\ell\nu$ suppression requirement rejects 98% of $B\rightarrow X_{c}\ell\nu$ events while retaining 25% of $B\rightarrow X_{u}\ell\nu$ events.
Figure 2: $B\rightarrow X_{c}\ell\nu$ suppression classifier output score. Simulated templates are shown as stacked histograms and experimental data are shown as black points. The signal component is shown as a purple line with its yield multiplied by 10 to enhance visibility. The bottom panel shows the ratio between data and simulation yields in each bin. The purple arrows indicate that the point is outside of the plotting range and the purple crosses indicate the absence of simulated data in a particular bin. The step in the distribution in the range 0.700.750.70-0.75 is caused by the large number of events with a $p_{\rm vtx}^{\rm ROE}$ close to zero (see Fig. 10 in Appendix A).
@@ IV.1 Continuum modeling corrections
The modeling of continuum events is known to be poorly understood and is therefore corrected using a data-driven approach. First, the total number of expected events is corrected by comparing the number of events in 40% of the off-resonance experimental and simulated data samples. The remaining 60% are used to test and validate the BDT described below. The correction factor is extracted independently for events in which a charged or neutral $B$ meson candidate has been reconstructed and is found to be $0.87\pm 0.02$ and $0.94\pm 0.03$, respectively, where the uncertainties are statistical (uncertainties on these factors are derived as described in Sec. V.5.9).
In a second step, a BDT with a maximum tree depth of five is trained to distinguish between experimental and simulated off-resonance data using 30 variables that describe the event shape [49] (see Sec. III.3). In addition, the missing mass squared and the scalar sum of the transverse momenta of the signal and tag-side $B$ final-state particles are included in the BDT input features. The BDT is trained on the off-resonance sample from which the normalization correction factors are extracted. The remainder of the sample is split into a validation sample (10% of the full sample) and a test sample (50%). The classifier output score $\mathcal{S}_{\rm cont}$ is well calibrated and takes values between 0 and 1. The ratio
@@
is an approximation to the ratio of experimental to simulated data likelihoods: it is applied as a weight to continuum events in the on-resonance simulated data sample. The continuum correction includes both the overall normalization factors and the BDT-extracted shape factors. The modeling correction is applied on the continuum sample after applying the preselections and before training the continuum suppression classifier. Its impact is illustrated in Fig. 3 for the most discriminating variable in the BDT training.
Figure 3: Distribution of $C_{1}$ (first CLEO cone [11]) in an off-resonance sample before (left) and after (right) the continuum modeling correction is applied. The modeling uncertainties, described in Sec. V.5.9, are included in the error band after applying the continuum calibration. The bottom panel shows the ratio between data and simulation yields in each bin. The purple arrows indicate that the point is outside of the plotting range and the purple crosses indicate the absence of simulated data in a particular bin.
@@ IV.2 $B\rightarrow X_{c}\ell\nu$ modeling corrections
The agreement between data and simulation is evaluated in control regions. Three control regions are defined for each kinematic selection defined in Table 2 by inverting the kaon veto and/or requiring a low $B\rightarrow X_{c}\ell\nu$ suppression classifier score (see Fig. 2):
@@
A comparison of the $B\rightarrow X_{c}\ell\nu$ composition in the three control regions and the signal region is shown in Appendix B. In addition, two validation regions, VRK,mid and VR0,mid, with classifier scores between 0.60 and 0.87 are defined. In the former region, a kaon is required whereas in the latter, kaons are vetoed. The subdivided selection plane is illustrated in Fig. 4. In Fig. 5, clear disagreements between data and simulation distributions of $E_{\ell}^{B}$, $M_{X}$ and $q^{2}$ can be seen.
@@
Background originating from continuum as well as fake and secondary leptons is known to be poorly modeled in simulation. However, continuum events are corrected following the procedure described in Sec. IV.1 and these background sources represent in total less than 10% of all events in the control regions. The mismodeling is therefore expected to stem from $B\rightarrow X_{c}\ell\nu$ events. There are several known shortcomings in the simulation of $B\rightarrow X_{c}\ell\nu$ decays. The $B\to D^{(*)}\eta\ell\nu$ modes have never been observed and are used as a hypothesis for the decays that fill the $B\rightarrow X_{c}\ell\nu$ gap. Therefore, a relative uncertainty of 100% is assigned to their assumed branching fractions. This uncertainty doesn’t cover the full difference between data and simulation distributions shown in Fig. 5. Furthermore, most $D$ meson decays with three or more particles are not simulated according to measured decay distributions but use simple phase-space models, which can bias the kinematics of charmed $B$ decays. Moreover, it was observed in Belle II studies [2] that measurements can be affected by the poor modeling of charm decays involving $K^{0}_{L}$ mesons. The impact of the modeling of $B\rightarrow X_{c}\ell\nu$ events with a $D$ meson decaying to a $K^{0}_{L}$ is studied in detail by scaling their contribution by arbitrary factors. Based on studies performed in the measurement of the $B^{+}\to K^{+}\nu\bar{\nu}$ branching fraction [2], this component is increased by 30% which only marginally improves the agreement between data and simulation. This indicates that it is unlikely that the mismodeling of $D\to K^{0}_{L}X$ decays can explain the observed mismodeling. As no single source of mismodeling could be identified, the $B\rightarrow X_{c}\ell\nu$ component is corrected as a whole using the control regions defined above. The normalization and shape from simulation are corrected separately.
Figure 5: $E_{\ell}^{B}$, $M_{X}$ and $q^{2}$ distributions in data and simulation in the CR0,low region defined in the text. Only the $E_{\ell}^{B}>1$ GeV kinematic selection is applied. All systematic uncertainties as described in Sec. V.5 are included. The bottom panel shows the ratio between data and simulation yields in each bin. The purple arrows indicate that the point is outside of the plotting range.
@@ IV.3 $B\rightarrow X_{c}\ell\nu$ normalization correction
The $B\rightarrow X_{c}\ell\nu$ normalization correction factors $r_{0,\mathrm{low}}$, $r_{K,\mathrm{low}}$ and $r_{K,\mathrm{high}}$ are calculated by subtracting from data and simulation all $B\rightarrow X_{u}\ell\nu$, continuum, as well as fake and secondary lepton simulated events. These components represent between 2% and 10% of all events in CR0,low and CRK,low and between 8% and 25% of events in CRK,high, depending on kinematic selections. Since this computation relies on various sources of background and signal being well modeled, an uncertainty is derived on the three ratios by varying the total yields of $B\rightarrow X_{u}\ell\nu$, continuum and fake and secondary lepton events by 50% and computing the normalization correction factors for each variation. The calculation defined in Eq. 9 is tested by replacing CRK,high and the signal region by VRK,mid and VR0,mid  respectively. Although the signal purity is larger in VR0,mid (about 10–20% depending on kinematic selections) than in all other control regions, it remains lower than in the signal region (30–70%). The $B\rightarrow X_{c}\ell\nu$ normalization correction obtained for VR0,mid via the control region method (1.07, 1.15, 1.21, from the broadest set of kinematic selections to the tightest) and the factor obtained by directly taking the $B\rightarrow X_{c}\ell\nu$ data over simulated data yield ratio in that region (1.03, 1.16, 1.22) are compared. The relative difference between these two factors is considered as an additional uncertainty on $r_{\rm SR}$. In the signal extraction, two nuisance parameters are included to float the normalization of the $B\rightarrow X_{c}\ell\nu$ component in CR0,low and in the signal region separately. The normalization parameters are constrained by Gaussian functions with widths equal to the uncertainty on the normalization correction of each region. The values of $r_{0,\mathrm{low}}$ and $r_{\rm SR}$ in the three phase-space regions considered in the signal extraction are given in Table 3.
@@
The $B\rightarrow X_{c}\ell\nu$ shape correction is implemented directly in the signal extraction procedure by simultaneously fitting the signal region and CR0,low as described in Sec. V.2.